# Validity of the Fröhlich model for a mobile impurity in a Bose-Einstein condensate

### v2

@@ I Introduction
The concept of a polaron, a quasi-particle formed through the interaction of an impurity with a medium, is ubiquitous in condensed matter physics [1, 2, 3]. Among the many examples, the Bose polaron stands out by the degree of control over the interactions allowed by cold atomic gases. For this reason, it is seen as a possible platform for the simulation of polaron physics in analogous systems, and in particular allows to access the large-coupling regime [4, 5, 6]. The system also serves a similar role for theory, as a many-body quantum system whose constituents are simple enough that one can understand its complex behavior and effective degrees of freedom from first principles and has thus attracted considerable attention [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. For the same reason, the rigorous justification of the effective description of fluctuations in dilute Bose gases by Bogoliubov phonons is an active area of mathematical research [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], with recent results also considering impurities [37, 38, 39].
Polarons are often modeled by the Fröhlich Hamiltonian, in which the impurity interacts with the field of low-energy excitations of the medium by linear coupling. However, the validity of this model for the Bose polaron has been questioned [11]. Given the importance of the Bose polaron as a model system, it is crucial to understand precisely the regime of validity of the Fröhlich Hamiltonian and control the modelling errors. In this letter we give conditions under which the Fröhlich Hamiltonian provides a good approximation to the full many-body spectrum of a dilute Bose gas with an impurity, and provide rigorous error estimates. This allows the model to be used with a great deal of confidence, and certifies the analogies to systems with a similar Hamiltonian.
@@
is of order one (definitions in [7, 2] differ by a factor of $8\pi$). That is, the impurity-boson interaction has considerably longer range than the boson-boson interaction. In experiments probing the strong-coupling regime of the Bose polaron [4, 5, 6], this corresponds to moderate interactions.
@@ II The model
We consider a single impurity interacting with $N$ bosons in a box of size $L=\xi$ with periodic boundary conditions at zero temperature. In units where $\hbar=\xi=1$, and also the impurity mass is $M=1$, the Hamiltonian is
$\hat{H}=\tfrac{1}{2}\hat{\bm{P}}^{2}+\sum_{j=1}^{N}\frac{\hat{\bm{p}}_{j}^{2}}{2m}+\sum_{1\leq i<j\leq N}V_{\mathrm{BB}}({\bm{x}}_{i}-{\bm{x}}_{j})+\sum_{j=1}^{N}V_{\mathrm{IB}}({\bm{X}}-{\bm{x}}_{j}),$
@@
[…]
@@
where $\mu^{-1}=1+m^{-1}$ is the reduced mass, $b_{\bm{k}}^{\dagger}$, $b_{{\bm{k}}}$ create and annihilate Bogoliubov phonons, whose dispersion relation is
@@
II.1 Asymptotic energy
Theorem 1.
Let $\tilde{V}_{\mathrm{BB}}$, $\tilde{V}_{\mathrm{IB}}$ be positive, have finite support and be square-integrable, and assume Bose–Einstein condensation occurs. Then, as $N=\xi/(8\pi a_{\mathrm{BB}})\to\infty$ the ground state energy in units of $1/\xi^{2}$ is
$E_{0}$ $=\frac{1}{4m}N+\frac{2\pi}{\mu}\sqrt{\alpha}N^{1/2}$
@@ II.1 Asymptotic energy
Moreover, the excited eigenvalues satisfy
$E_{n}-E_{0}=E_{n}^{\mathrm{BF}}-E_{0}^{\mathrm{BF}}+\mathcal{O}(N^{-\delta})$
for some $\delta>0$.
In this expansion, the leading term corresponds to the usual Gross-Pitaevskii energy $4\pi a_{\mathrm{BB}}\varrho_{B}^{2}L^{3}$. The following term is the analogue for the boson-impurity interaction $8\pi a_{\mathrm{IB}}\varrho_{\mathrm{B}}\varrho_{\mathrm{I}}L^{3}$. The term of order $\log N$ arises from an effective phonon-phonon interaction mediated by the impurity, and thus vanishes in the limit of a heavy impurity, $M\to\infty$. Such a term has been observed in the context of the Fröhlich-Hamiltonian [10, 45], and our study shows that it is the leading correction to the mean-field approximation. Similar terms can be found in the expansion of the energy in $a_{\mathrm{BB}}/\xi$ (without impurity) [40, 41, 42, 43], and in $a_{\mathrm{IB}}/\xi$ [11], but in our parameter regime these are of lower order ($N^{-1}\log N$, resp. $N^{-1/2}\log N$). This energy shift may thus be much easier to detect experimentally than the corresponding terms in the pure Bose gas, which so far is only theoretical. Indeed, the shift is even larger than the Lee-Huang-Yang [46] term, which is of order one in our units and has been observed [47, 48, 49, 50].
@@ II.2 The 2+1–body problem
In relative coordinates ${\bm{r}}_{i}={\bm{X}}-{\bm{x}}_{i}$, $i=1,2$, the Hamiltonian with $V_{\mathrm{BB}}=0$ is
$\hat{H}_{2+1}=-\tfrac{1}{2\mu}(\Delta_{{\bm{r}}_{1}}+\Delta_{{\bm{r}}_{2}})+\nabla_{{\bm{r}}_{1}}\cdot\nabla_{{\bm{r}}_{2}}+V_{\mathrm{IB}}({\bm{r}}_{1})+V_{\mathrm{IB}}({\bm{r}}_{2}).$
Given the short range $a_{\mathrm{IB}}=N^{-1/2}\sqrt{\alpha}$ of the interaction, two-body effects will dominate and we make the ansatz $\Psi({\bm{r}}_{1},{\bm{r}}_{2})=f({\bm{r}}_{1})f({\bm{r}}_{2})$ for a low-energy state, with $f$ normalized, periodic. Calculating the energy yields
$\langle\Psi,\hat{H}_{2+1}\Psi\rangle=-\frac{1}{\mu}\int\overline{f({\bm{r}})}\Delta f({\bm{r}})+2\int V_{\mathrm{IB}}({\bm{r}})f({\bm{r}})^{2}d^{3}r,$ (4)
as the term with $\nabla_{{\bm{r}}_{1}}\cdot\nabla_{{\bm{r}}_{2}}$ integrates to zero. The minimizer $f$ of the right hand side solves a periodized version of the zero-energy scattering equation ($f\propto 1+\varphi_{\mathrm{IB}}$ from (9)). Inserting this into (4) gives the leading contribution to the energy $2E_{\text{2-bd}}\sim 4\pi N^{-1/2}\sqrt{\alpha}/\mu$. In order to obtain a more precise result, we refine the ansatz by taking into account 3-body effects, $\Psi({\bm{r}}_{1},{\bm{r}}_{2})=f({\bm{r}}_{1})f({\bm{r}}_{2})\Phi({\bm{r}}_{1},{\bm{r}}_{2})$, where $\Phi$ is normalized and differs significantly from one only where $r_{1},r_{2}\lesssim a_{\mathrm{IB}}$. Then, the mixed derivatives no longer integrate to zero and play the role of an effective potential
@@
Minimizing in $\Phi$, the three-body contribution is a generalized scattering energy of $V_{\mathrm{eff}}$ [51]. It evaluates approximately to
@@
$\approx 32\pi\alpha^{2}\mu^{-1}\Big{(}\mu^{-1}\arcsin\mu-\sqrt{1-\mu^{2}}\Big{)}N^{-2}\log N.$
The term in the expansion of the $N$-body energy is obtained by multiplying $E_{\text{3-bd}}$ by $N^{2}/2$, or the number of boson pairs.
@@ III Derivation of the Fröhlich Hamiltonian
We start by rewriting the Hamiltonian in second quantization and separate the condensate particles with ${\bm{k}}=0$ from the excitations with ${\bm{k}}\neq 0$, which yields
$\hat{H}=\tfrac{1}{2}N_{0}(N_{0}-1)\hat{V}_{\mathrm{BB}}(0)+N_{0}\hat{V}_{\mathrm{IB}}(0)$ (6a)
$\quad+\tfrac{1}{2}\hat{P}^{2}+\frac{1}{2m}\sum_{{\bm{k}}\neq 0}k^{2}$ (6b)
@@
$\quad+\frac{1}{2}\sqrt{N_{0}}\sum_{{\bm{k}},{\bm{\ell}}\neq 0}\hat{V}_{\mathrm{BB}}({\bm{k}})a^{\dagger}_{{\bm{k}}+{\bm{\ell}}}a_{\bm{k}}a_{\bm{\ell}}+\mathrm{h.c.}$ (6g)
$\quad+\frac{1}{2}\sum_{{\bm{k}},{\bm{\ell}},\bm{\neq}0}\hat{V}_{\mathrm{BB}}({\bm{k}})a^{\dagger}_{{\bm{k}}+{\bm{\ell}}}a^{\dagger}_{{\bm{\ell}}}a_{\bm{m}}a_{\bm{m}-{\bm{k}}}$ (6h)
@@
Our condensation hypothesis on $\tilde{V}_{\mathrm{BB}}$ means that on low-energy states $N_{0}\geq N-c\sqrt{N}$. In view of the scaling relations (1)–(2), we are thus tempted to neglect the terms (6f)–(6h). We would then arrive at a Fröhlich-type Hamiltonian and an energy expansion similar to Theorem 1. However, these would feature $\hat{V}_{\mathrm{BB}}(0)$, $\hat{V}_{\mathrm{IB}}(0)$ instead of the scattering lengths $a_{\mathrm{BB}}$, $a_{\mathrm{IB}}$ and thus be incorrect even at the leading order. Because interactions in $\hat{H}$ are strong on short length scales, particles correlate to balance the kinetic energy with the interaction. As a result, although most modes have a momentum of order one (that is, $\xi^{-1}$ in our units), a vanishing number of them have large momenta and are responsible for a shift in energy to leading order. We first need to factor out the short scale correlation structure, making the scattering length appear in the couplings instead of the bare potentials and changing the kinetic energy from order $N$ to order one on low energy states.
After this first step, we are left with quadratic and linear terms acting on low momenta, which up to a scalar contributions is essentially an ultraviolet cutoff version of
@@ III.1 Approximate diagonalization
Large momenta are those with $k\gtrsim N^{\tau}$, for some small $20<\tau<1/$ 2. The unitary transformations are obtained by exponentiating linear, quadratic and cubic operators
@@
for ${\bm{k}}\neq 0$, and $\chi(N_{+})$ truncates the excitation number.
@@
From the equation (8), we see that $\varphi_{\mathrm{BB}}({\bm{k}})$ behaves similarly to $-\hat{V}_{\mathrm{BB}}({\bm{k}})/(mk^{2})$, and thus $\sum_{k>N^{\tau}}|N\varphi_{\mathrm{BB}}({\bm{k}})|^{2}\sim N^{-\tau}$. The unitaries are therefore close to the identity. However, since, e.g., $\sum_{k>N^{\tau}}k^{2}|N\varphi_{\mathrm{BB}}({\bm{k}})|^{2}\sim N$, they affect the energy to leading order.
@@
[…]
The renormalization of the energy and the couplings becomes apparent when evaluating these commutators and simplifying the result using the scattering equations (8), (9). The error terms are small relative to the kinetic operators in (6c) and the positive interaction terms (6f), (6h).
@@
This transformation implements the correlations due to hard boson pairs. It acts non-trivially on the boson Hamiltonian, namely on the terms (6b),(6d) and (6h). Using the scattering equation (8) leads to the replacements
@@
Now that the kinetic energy is of order $N^{\tau}$, we can perform the cubic transformation, whose role is to take into account correlations of soft boson pairs. Acting mainly on the terms (6b),(6g) and (6h), it factors out contributions from the cubic term and renormalizes the diagonal quadratic terms of the boson-boson interaction

### v3

@@ I Introduction
The concept of a polaron, a quasi-particle formed through the interaction of an impurity with a medium, is ubiquitous in condensed matter physics [1, 2, 3]. Among the many examples, the Bose polaron stands out by the degree of control over the interactions allowed by cold atomic gases. For this reason, it is seen as a possible platform for the simulation of polaron physics in analogous systems, and in particular allows to access the large-coupling regime [4, 5, 6]. The system also serves a similar role for theory, as a many-body quantum system whose constituents are simple enough that one can understand its complex behavior and effective degrees of freedom from first principles and has thus attracted considerable attention [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. For the same reason, the rigorous justification of the effective description of fluctuations in dilute Bose gases by Bogoliubov phonons is an active area of mathematical research [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], with recent results also considering impurities [37, 38, 39].
Polarons are often modeled by the Fröhlich Hamiltonian, in which the impurity interacts with the field of low-energy excitations of the medium by linear coupling. However, the validity of this model for the Bose polaron has been questioned [11]. Given the importance of the Bose polaron as a model system, it is crucial to understand precisely the regime of validity of the Fröhlich Hamiltonian and control the modeling errors. In this paper, we give conditions under which the Fröhlich Hamiltonian provides a good approximation to the full many-body spectrum of a dilute Bose gas with an impurity. Our arguments are based on the mathematical analysis of [39] where we prove rigorous error estimates. This allows the model to be used with a great deal of confidence, and certifies the analogies to systems with a similar Hamiltonian.
@@
is of order one  [^1] Note that defintions of this parameter in [7, 2] differ by a factor of $8\pi$) . That is, the impurity-boson scattering length is considerably larger than that of the boson-boson interaction. In experiments probing the strong-coupling regime of the Bose polaron [4, 5, 6], this corresponds to moderate interactions.
@@ II The model
We consider a single impurity interacting with $N$ bosons in a box of size $L=\xi$, with periodic boundary conditions at zero temperature. That is, $N$ is proportional to $(\varrho_{\mathrm{B}}a_{\mathrm{BB}}^{3})^{-1/2}$, which is sometimes called the Gross–Pitaevskii regime. In units where $\hbar=\xi=1$, and also the impurity mass is $M=1$, the Hamiltonian is
$\hat{H}=\tfrac{1}{2}\hat{\bm{P}}^{2}+\sum_{j=1}^{N}\frac{\hat{\bm{p}}_{j}^{2}}{2m}+\hskip-4.0pt\sum_{1\leq i<j\leq N}\hskip-9.0ptV_{\mathrm{BB}}({\bm{x}}_{i}-{\bm{x}}_{j})+\sum_{j=1}^{N}V_{\mathrm{IB}}({\bm{X}}-{\bm{x}}_{j}),$
@@
where now $\tilde{a}_{\mathrm{BB}}=Na_{\mathrm{BB}}=1/8\pi$ (cf. the definition  [^2] We use here the definition of the scattering length of a general potential $4\pi a=\min\int|\nabla\varphi(x)|^{2}+2m_{\mathrm{red}}(1+\varphi(x))^{2}V(x)dx=2m_{\mathrm{red}}\int V(x)(1+\varphi_{\mathrm{min}}(x))dx$). Similarly, $a_{\mathrm{IB}}/\xi=\sqrt{\alpha}N^{-1/2}$ and we write
@@
where $\mu^{-1}=1+m^{-1}$ is the inverse reduced mass, $b_{\bm{k}}^{\dagger}$, $b_{{\bm{k}}}$ create and annihilate Bogoliubov phonons, whose dispersion relation is
@@
Here, and in the following, sums over momenta are over ${\bm{k}}\in 2\pi\mathbb{Z}/\xi$.
@@
II.1 Energy asymptotics
Our main results concern the asymptotic expansion of the energy for $N=\xi/(8\pi a_{\mathrm{BB}})\gg 1$ for repulsive, finite range interactions $\tilde{V}_{\mathrm{BB}}$, $\tilde{V}_{\mathrm{IB}}\geq 0$ (see [39, Theorem 1.2] for precise assumptions). The first is an expansion of the ground state energy of $\hat{H}$ in units of $1/\xi^{2}$
$E_{0}$ $=\frac{1}{4m}N+\frac{2\pi}{\mu}\sqrt{\alpha}N^{1/2}$ (4)
@@ II.1 Energy asymptotics
In this expansion, the leading term corresponds to the usual Gross-Pitaevskii energy $4\pi a_{\mathrm{BB}}\varrho_{B}^{2}L^{3}$. The following term is the analogue for the boson-impurity interaction $8\pi a_{\mathrm{IB}}\varrho_{\mathrm{B}}\varrho_{\mathrm{I}}L^{3}$. The term of order $\log N$ arises from an effective phonon-phonon interaction mediated by the impurity. It vanishes in the limit of a heavy impurity, that is, if instead of taking $M=1$ we let $M\to\infty$. Such a term has been observed in the context of the Fröhlich-Hamiltonian [10, 46], and our study shows that it is the leading correction to the mean-field approximation. Terms that arise from similar effective three-body interactions can be found in the expansion of the energy in $a_{\mathrm{BB}}/\xi$ (without impurity) [40, 41, 42, 43], and in $a_{\mathrm{IB}}/\xi$ [11], but in our parameter regime these are of lower order ($N^{-1}\log N$, resp. $N^{-1/2}\log N$). This energy shift may thus be much easier to detect experimentally than the corresponding term in the pure Bose gas, which so far is only theoretical. Indeed, the shift is even larger than the Lee-Huang-Yang [47] term, which is of order one in our units and has been observed [48, 49, 50, 51].
Our second result relates the excited eigenvalues $E_{n}$ to the corresponding eigenvalues $E_{n}^{\mathrm{BF}}$ of the Bogoliubov-Fröhlich Hamiltonian $\hat{H}_{\mathrm{BF}}$. We have
$E_{n}-E_{0}=E_{n}^{\mathrm{BF}}-E_{0}^{\mathrm{BF}}+\mathcal{O}(N^{-\delta})$ (5)
for some $\delta>0$. This shows that the excitation spectrum is well approximated by that of $\hat{H}_{\mathrm{BF}}$, and thus justifies the latter as an effective description. Our analysis also yields an approximation of the eigenstates and the dynamics, see [39, Corollary 1.6].
@@ II.2 The 2+1–body problem
Ignoring the boson-boson interaction, we consider the following Hamiltonian
$\hat{H}_{2+1}=-\tfrac{1}{2\mu}(\Delta_{{\bm{r}}_{1}}+\Delta_{{\bm{r}}_{2}})+\nabla_{{\bm{r}}_{1}}\cdot\nabla_{{\bm{r}}_{2}}+V_{\mathrm{IB}}({\bm{r}}_{1})+V_{\mathrm{IB}}({\bm{r}}_{2}),$
where ${\bm{r}}_{i}$, $i=1,2$, denote the relative coordinates between each boson and the impurity. Due to the diluteness of the system, two-body effects will dominate and we make the ansatz $\Psi({\bm{r}}_{1},{\bm{r}}_{2})=f({\bm{r}}_{1})f({\bm{r}}_{2})$ for a low-energy state, with $f$ normalized, periodic. Calculating the energy yields
$\langle\Psi,\hat{H}_{2+1}\Psi\rangle=-\frac{1}{\mu}\int\overline{f({\bm{r}})}\Delta f({\bm{r}})d^{3}r+2\int V_{\mathrm{IB}}({\bm{r}})f({\bm{r}})^{2}d^{3}r,$ (6)
as the term with $\nabla_{{\bm{r}}_{1}}\cdot\nabla_{{\bm{r}}_{2}}$ integrates to zero. The minimizer $f$ of the right hand side solves a periodized version of the zero-energy scattering equation ($f\propto 1+\varphi_{\mathrm{IB}}$ from (11)). Inserting this into (6) gives the leading contribution to the energy $2E_{\text{2-bd}}\sim 4\pi N^{-1/2}\sqrt{\alpha}/\mu$. In order to obtain a more precise result, we refine the ansatz by taking into account 3-body effects, $\Psi({\bm{r}}_{1},{\bm{r}}_{2})=f({\bm{r}}_{1})f({\bm{r}}_{2})\Phi({\bm{r}}_{1},{\bm{r}}_{2})$, where $\Phi$ differs significantly from one only where $r_{1},r_{2}\lesssim a_{\mathrm{IB}}$. Then, the mixed derivatives no longer integrate to zero and play the role of an effective potential
@@
Minimizing in $\Phi$, the three-body contribution is a generalized scattering energy of $V_{\mathrm{eff}}$ [52]. Second order peturbation theory gives
@@
$\approx 32\pi\alpha^{2}\mu^{-1}\Big(\mu^{-1}\arcsin\mu-\sqrt{1-\mu^{2}}\Big)N^{-2}\log N,$
where the second approximation replaces the sum by an integral, evaluated as in [39, Eq. (6.45)]. The term in the expansion of the $N$-body energy is obtained by multiplying $E_{\text{3-bd}}$ by $N^{2}/2$, or the number of boson pairs.
@@ III Derivation of the Fröhlich Hamiltonian
We start by rewriting the Hamiltonian in second quantization and using the isomorphism of [28], to separate the condensate particles with ${\bm{k}}=0$ from the excitations with ${\bm{k}}\neq 0$ and work on the truncated Fock space of excitations $\mathcal{F}^{\leq N}\left(\ell^{2}(2\pi\mathbb{Z}^{3}/\{0\}\right))$:
$\hat{H}\simeq\tfrac{1}{2}N_{0}(N_{0}-1)\hat{V}_{\mathrm{BB}}(0)+N_{0}\hat{V}_{\mathrm{IB}}(0)$ (8a)
$\quad+\tfrac{1}{2}\hat{P}^{2}+\frac{1}{2m}\sum_{{\bm{k}}\neq 0}k^{2}a_{\bm{k}}^{\dagger}a_{\bm{k}}$ (8b)
@@
$\quad+\frac{1}{2}\sqrt{N_{0}}\sum_{{\bm{k}},{\bm{\ell}},{\bm{k}}+{\bm{\ell}}\neq 0}\hat{V}_{\mathrm{BB}}({\bm{k}})a^{\dagger}_{{\bm{k}}+{\bm{\ell}}}a_{\bm{k}}a_{\bm{\ell}}+\mathrm{h.c.}$ (8g)
$\quad+\frac{1}{2}\sum_{{\bm{\ell}}+{\bm{k}},\bm{m}\neq 0\atop\bm{m}+{\bm{k}},{\bm{\ell}}\neq 0}\hat{V}_{\mathrm{BB}}({\bm{k}})a^{\dagger}_{{\bm{\ell}}+{\bm{k}}}a^{\dagger}_{\bm{m}}a_{\bm{m}+{\bm{k}}}a_{{\bm{\ell}}}$ (8h)
@@
Low energy states of this Hamiltonian exhibit complete Bose–Einstein condensation, $N_{0}\approx N$. In view of the scaling relations (1)–(2), we are thus tempted to neglect the terms (8f)–(8h). We would then arrive at a Fröhlich-type Hamiltonian and an energy expansion as described above. However, these would feature $\hat{V}_{\mathrm{BB}}(0)$, $\hat{V}_{\mathrm{IB}}(0)$ instead of the scattering lengths $a_{\mathrm{BB}}$, $a_{\mathrm{IB}}$ and thus be incorrect even at the leading order. Because interactions in $\hat{H}$ are strong on short length scales, particles correlate to balance the kinetic energy with the interaction. As a result, although most modes have a momentum of order one (that is, $\xi^{-1}$ in our units), a vanishing number of them have large momenta and are responsible for a shift in energy to leading order. We first need to factor out the short scale correlation structure, making the scattering length appear in the couplings instead of the bare potentials and changing the kinetic energy from order $N$ to order one on low energy states.
After this first step, we are left with quadratic and linear terms acting on low momenta, which up to an additive constant is essentially an ultraviolet cutoff version of
@@ III.1 Approximate diagonalization
Large momenta are those with $k\gtrsim N^{\tau}$, for some small $0<\tau<1/2$. The unitary transformations are obtained by exponentiating linear, quadratic and cubic operators
@@
for ${\bm{k}}\neq 0$, and $\chi(N_{+})$ restricts the excitation number $N_{+}=N-N_{0}$ to values much smaller than $N$.
@@
From the equation (10), we see that $\varphi_{\mathrm{BB}}({\bm{k}})$ behaves similarly to $m\hat{V}_{\mathrm{BB}}({\bm{k}})/k^{2}$, and thus, using (1), we obtain $\sum_{k>N^{\tau}}|N\varphi_{\mathrm{BB}}({\bm{k}})|^{2}\sim N^{-\tau}$. The unitaries are therefore close to the identity. However, since, e.g., $\sum_{k>N^{\tau}}k^{2}|N\varphi_{\mathrm{BB}}({\bm{k}})|^{2}\sim N$, they affect the energy to leading order.
@@
$e^{-B}Ae^{B}$ $=A+[A,B]+\int_{0}^{1}du\int_{0}^{u}dte^{-tB}[[A,B],B]e^{tB}.$
The renormalization of the energy and the couplings becomes apparent when evaluating these commutators and simplifying the result using the scattering equations (10), (11). The error terms are small relative to the kinetic operators in (8b) and the positive interaction terms (8f), (8h).
@@
This transformation implements the correlations associated to hard boson pairs. It acts non-trivially on the boson Hamiltonian, namely on the terms (8b),(8d) and (8h). We mean by this that the other terms remain unchanged up to small errors relative to the desired precision. Using the scattering equation (10) leads to the replacements
@@
Now that the kinetic energy is of order $N^{\tau}$ on low energy states, we can perform the cubic transformation, whose role is to take into account correlations of soft boson pairs. Acting mainly on the terms (8b),(8g) and (8h), it factors out contributions from the cubic term and renormalizes the diagonal quadratic terms of the boson-boson interaction